OpenStudy (ayeeeitskateee):

Need help with elimination. I would like somebody to please help me work through the problem. I do like to write them down so please allow at least 30 seconds for me to write what you said. It would be greatly appreciated if somebody would help me LEARN. No answers until I figure it out for myself, please! Thanks! Equation below. 2x+6y=-12 5x-5y=10. I know I have to multiply by a number for the first step but that's the problem, I don't know what number. How do I know? Answer this question first please. After that though, please help me work through it. :)) Thanks guys! Will medal/fan

OpenStudy (ayeeeitskateee):

Greatly appreciate any help, even tips! thanks all.

OpenStudy (dayakar):

2x+6y=-12 divide each term with 2,u will get x+3y = -6 --------I sx-5y=10 divide each term with 5 x-y= 2--------II

OpenStudy (dayakar):

subtract II from I

OpenStudy (dayakar):

4y= -8 y= -8/4 y=-2

OpenStudy (ayeeeitskateee):

Oh. so we have to divide? I thought we had to multiply. Let me read the rest of your response. :)

OpenStudy (dayakar):

it can be done by multiplication also but you have to deal with some big numbers

OpenStudy (ayeeeitskateee):

I just wrote it down and i understand now. Thank you very much. :)) Will medal and fan be okay? :)

OpenStudy (ayeeeitskateee):

Oh. Okay. Well, how about x now that we found y?

OpenStudy (ayeeeitskateee):

Plug it in maybe?

OpenStudy (dayakar):

ok substitute y=-2 in one of eqations u get x value can u do that

OpenStudy (ayeeeitskateee):

Mmhm. I will start writing now. What equation do I plug it in? Or doesn't it matter?

OpenStudy (ayeeeitskateee):

Ok, I got ya! Didn't read it correctly the first time. So it is either equation.

OpenStudy (dayakar):

right

OpenStudy (ayeeeitskateee):

0?

OpenStudy (dayakar):

excellent

OpenStudy (ayeeeitskateee):

So, (0,-2)?

OpenStudy (ayeeeitskateee):

Thank you so much!!

OpenStudy (dayakar):

right

OpenStudy (ayeeeitskateee):

One more thing.

OpenStudy (dayakar):

ok,any more

OpenStudy (ayeeeitskateee):

What is the solution of the following system? -3x-2y=-12 9x+6y=-9. How do I know if i should substitute or eliminate? That's all I need to know for this. :)

OpenStudy (dayakar):

in first eqation multiply with (-) both sides

OpenStudy (dayakar):

3x+2y=12-------I

OpenStudy (ayeeeitskateee):

Okay, so what would that be? Substitution or elimination?

OpenStudy (dayakar):

second equation divide with 3 both sides

OpenStudy (dayakar):

3x+2y=-3-------II

OpenStudy (ayeeeitskateee):

okay, I am writing. :)

OpenStudy (ayeeeitskateee):

Got it. Then what?

OpenStudy (ayeeeitskateee):

Hm. They look almost the same.

OpenStudy (dayakar):

in both eqations has x,y coefficients equal

OpenStudy (dayakar):

except constant term, here RHS PART

OpenStudy (ayeeeitskateee):

What does that mean? RHS PART?

OpenStudy (dayakar):

THEREFORE TWO EQUATIONS SHOWS parellel lines

OpenStudy (dayakar):

right hand side=RHS

OpenStudy (ayeeeitskateee):

Does that mean many solutions? That is a choice in my book.

OpenStudy (ayeeeitskateee):

Sorry if I'm wrong.

OpenStudy (dayakar):

NO SOLUTION

OpenStudy (ayeeeitskateee):

Oh. Okay. Why would it be no solution?

OpenStudy (dayakar):

BECAUSE,two lines never intersect,how can u get common point mean solution

OpenStudy (ayeeeitskateee):

Ohhh. I got it now! Thank you. You are a life saver. :) Thank you for the knowledge. I love learning. That is all I need. :)

OpenStudy (dayakar):

wait

OpenStudy (ayeeeitskateee):

hm?

OpenStudy (dayakar):



OpenStudy (dayakar):

ok

OpenStudy (ayeeeitskateee):

I will save that for future reference. :)